# Pronunciation Editing for Finnish Speech using Phonetic Posteriorgrams

### v2

@@ 4.3.2 Subjective Evaluation
The results for subjective evaluation are shown in Table 5. Our Mean Opinion Score (MOS) tests have 16 native Finnish participants for each test and 75 utterances for each system in each test. We use language proficiency questions that instruct the participants to rate certain questions with specific scores in Finnish text and audio prompts to filter out invalid submissions. Our MOS tests include naturalness MOS (NMOS row), speaker similarity MOS (SMOS), and a pronunciation editing MOS test (EMOS rows). For NMOS, the participants are asked to evaluate whether the speech is natural. For SMOS, the participants are presented with the natural reference utterances and synthetic speech with different content, and they are asked to rate how similar the speakers of these two speech samples are. The synthetic speech is synthesized using the same speaker embedding from the reference groundtruth. For both NMOS and SMOS, the scores are between 1-5 with an increment of 1. For EMOS, we provide the corresponding transcripts after the PPG editing, and the edited speech to the participants, while they are required to answer a yes-or-no question on whether the edited speech agrees with the edited transcripts, and rate the naturalness of the edited speech the same as NMOS. Positive Rate (the ratio of the yes answer) is calculated as the subjective metric to the editing effect, which should be between 0 and 1. We use Student’s t-test to obtain the $95\%$ confidence interval of each MOS, and use the Wilcoxon signed-rank test to determine whether the differences in MOS of any pair of systems are significant.
@@
In Table 5, the Groundtruth has an NMOS of $4.27\pm 0.05$, indicating the moderate dataset quality we have for this study. Among speech synthesis models, PPG2Speech-CFG achieves the highest NMOS ($3.51\pm 0.07$). Moreover, comparing Matcha-TTS with Matcha-TTS-CFG using the Wilcoxon signed-rank test yields a p-value of $0.76$, meaning the difference on NMOS between Matcha-TTS and Matcha-TTS-CFG is insignificant, while the difference on NMOS between PPG2Speech and PPG2Speech-CFG is significant. This demonstrates that CFG can improve the naturalness in our experiments, and PPG2Speech-CFG can synthesize natural speech, and the slightly high CER shown in Table 3 doesn’t affect the listening experience. Matcha-TTS scores highest in SMOS ($3.63\pm 0.09$), closely followed by PPG2Speech ($3.57\pm 0.09$) and PPG2Speech-CFG ($3.54\pm 0.09$). The Wilcoxon signed-rank test between the SMOS on PPG2Speech and the SMOS on PPG2Speech-CFG results in a p-value of $0.61$, suggesting the marginal advantage achieved by PPG2Speech over PPG2Speech-CFG is insignificant, while the Wilcoxon signed-rank test results between other pairs of SMOS suggest a significant difference. The SMOS is consistent with the SECS in Table 3, suggesting that pitch sequences can improve naturalness yet bring timbre leak to the model. To our surprise, Matcha-TTS-CFG shows the best Positive Rate ($0.87\pm 0.02$) in the EMOS evaluation. We think this is caused by the undetermined nature of the PPG, which baffles the model from following the editing strictly. Meanwhile, PPG2Speech-CFG attains the highest NMOS within EMOS ($3.66\pm 0.06$).
@@ 5 Conclusions
In this paper, we present PPG2Speech, a diffusion-based multispeaker PPG-to-Speech model, capable of phoneme-level pronunciation editing to approximate L2 language learners. We strengthen the original Matcha-TTS flow-matching decoder with CFG. We also proposed a task-specific objective evaluation metric, Phonetic Aligned Consistency (PAC), to measure the effectiveness of phoneme-level editing. The objective and subjective evaluations indicate that our PPG2Speech model can synthesize natural speech in both synthesis and editing tasks. It also follows the speaker identity in the speaker embedding, although timbre leaks caused slight degradation in the SECS and SMOS. For the editing task, both methods achieve more than $80\%$ positive rate in the test samples. Our PPG2Speech model performs better than the text-based editing in PAC, yet is inferior to the text-based editing in subjective evaluation, indicating that PAC might not be able to reflect the actual editing effect in some cases. Our future work includes introducing foreign phonemes to the system and editing speech to mimic L2 speakers’ pronunciation errors on phoneme duration and misplacement of open and closed vowels.

### v3

@@ 4.3.2 Subjective Evaluation
The results for subjective evaluation are shown in Table 5. Our Mean Opinion Score (MOS) tests have 16 native Finnish participants for each test and 75 utterances for each system in each test. We use language proficiency questions that instruct the participants to rate certain questions with specific scores in Finnish text and audio prompts to filter out invalid submissions. Our MOS tests include naturalness MOS (NMOS row), speaker similarity MOS (SMOS), and a pronunciation editing MOS test (EMOS rows). For NMOS, the participants are asked to evaluate whether the speech is natural. For SMOS, the participants are presented with the natural reference utterances and synthetic speech with different content, and they are asked to rate how similar the speakers of these two speech samples are. The synthetic speech is synthesized using the same speaker embedding from the reference groundtruth. For both NMOS and SMOS, the scores are between 1-5 with an increment of 1. For EMOS, we provide the corresponding transcripts after the PPG editing, and the edited speech to the participants, while they are required to answer a yes-or-no question on whether the edited speech agrees with the edited transcripts, and rate the naturalness of the edited speech the same as NMOS. Positive Rate (the ratio of the yes answer) is calculated as the subjective metric to the editing effect, which should be between 0 and 1. We use Student's t-test to obtain the $95\%$ confidence interval of each MOS, and use the Wilcoxon signed-rank test to determine whether the differences in MOS of any pair of systems are significant.
@@
In Table 5, the Groundtruth has an NMOS of $4.27\pm 0.05$, indicating the moderate dataset quality we have for this study. Among speech synthesis models, PPG2Speech-CFG achieves the highest NMOS ($3.51\pm 0.07$). Moreover, comparing Matcha-TTS with Matcha-TTS-CFG using the Wilcoxon signed-rank test yields a p-value of $0.76$, meaning the difference on NMOS between Matcha-TTS and Matcha-TTS-CFG is insignificant, while the difference on NMOS between PPG2Speech and PPG2Speech-CFG is significant. This demonstrates that CFG can improve the naturalness in our experiments, and PPG2Speech-CFG can synthesize natural speech, and the slightly high CER shown in Table 3 doesn't affect the listening experience. Matcha-TTS scores highest in SMOS ($3.63\pm 0.09$), closely followed by PPG2Speech ($3.57\pm 0.09$) and PPG2Speech-CFG ($3.54\pm 0.09$). The Wilcoxon signed-rank test between the SMOS on PPG2Speech and the SMOS on PPG2Speech-CFG results in a p-value of $0.61$, suggesting the marginal advantage achieved by PPG2Speech over PPG2Speech-CFG is insignificant, while the Wilcoxon signed-rank test results between other pairs of SMOS suggest a significant difference. The SMOS is consistent with the SECS in Table 3, suggesting that pitch sequences can improve naturalness yet bring timbre leak to the model. To our surprise, Matcha-TTS-CFG shows the best Positive Rate ($0.87\pm 0.02$) in the EMOS evaluation. We think this is caused by the undetermined nature of the PPG, which baffles the model from following the editing strictly. Meanwhile, PPG2Speech-CFG attains the highest NMOS within EMOS ($3.66\pm 0.06$).
@@ 5 Conclusions
In this paper, we present PPG2Speech, a diffusion-based multispeaker PPG-to-Speech model, capable of phoneme-level pronunciation editing to approximate L2 language learners. We strengthen the original Matcha-TTS flow-matching decoder with CFG. We also proposed a task-specific objective evaluation metric, Phonetic Aligned Consistency (PAC), to measure the effectiveness of phoneme-level editing. The objective and subjective evaluations indicate that our PPG2Speech model can synthesize natural speech in both synthesis and editing tasks. It also follows the speaker identity in the speaker embedding, although timbre leaks caused slight degradation in the SECS and SMOS. For the editing task, both methods achieve more than $80\%$ positive rate in the test samples. Our PPG2Speech model performs better than the text-based editing in PAC, yet is inferior to the text-based editing in subjective evaluation, indicating that PAC might not be able to reflect the actual editing effect in some cases. Our future work includes introducing foreign phonemes to the system and editing speech to mimic L2 speakers' pronunciation errors on phoneme duration and misplacement of open and closed vowels.